# Linying Lv

### v2

@@ 3.2 Information Content of Analyst Reports
In this section, I examine the information content of analysts’ quantitative and qualitative outputs by projecting contemporaneous cumulative abnormal returns on analyst information. Furthermore, I attribute the information content of analyst report text across topics using the Shapley value decomposition introduced in Section 2.4.
3.2.1 Information Content of Quantitative and Qualitative Information
@@
Table 2 compares the information content embedded in numerical and textual measures, quantified with $R^{2}_{\mathrm{OOS}}$. The estimation model applied in this table is the Ridge regression specified in Equations 1 and 2. The objective variable is ${CAR}_{[-1,+1]}$ centered at the analyst report release date. For reports released on non-trading days, I adjust t=0 to the next trading day recorded in the CRSP database. I estimate separate Ridge regression models for distinct input configurations and employ an expanding window approach, constructing training sets from 2000 to the year preceding each target year.
@@
Panel A of the table delineates the performance of four types of input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
@@
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% with forecast revision input, corroborating extant literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst forecast revisions. The inclusion of additional numerical variables in column (2) yields only a marginal improvement, with the $R^{2}_{\mathrm{OOS}}$ reaching 9.08% using 17 numerical features. I compare the estimation with a zero benchmark and report the Diebold-Mariano test t-statistics in columns (2) and (4). The models with revision and numerical inputs significantly outperform a zero-estimation benchmark.
@@
Columns (5) and (6) elucidate the performance of Ridge models with text embeddings as input, showcasing a substantial $R^{2}_{\mathrm{OOS}}$ of 10.19% and a t-statistic of 8.20, underscoring the significant information content of the report text in explaining contemporaneous stock returns. The $R^{2}_{\mathrm{OOS}}$ exhibits relative stability across 2015-2023, with a notable exception during the 2020 pandemic recession. The DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, suggesting substantial information content in analyst report text compared with forecast revisions.
@@
Sarkar and Vafa (2024) raise the look-ahead bias concern associated with pre-trained LLMs. Language model outputs may incorporate future information that should not be available. Since the LLMs I applied were trained on language data ending before 2023, the $R^{2}_{\mathrm{OOS}}$ in 2023 reflects true out-of-sample performance, confirming the robustness of the main findings during this period.
@@
In columns (7) and (8), I further investigate the model performance when combining both forecast revisions and text embeddings in the Ridge regression. The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, which represents a 3.27% improvement over the $R^{2}_{\mathrm{OOS}}$ of the revision-only model and a 2.09% improvement over the $R^{2}_{\mathrm{OOS}}$ of the text-only model. I report the DM t-statistic of the comparison between combined input with revision input and text input in columns (7)-(1) and (7)-(5). The overall t-statistics of 3.95 and 3.77 indicate a significant superiority in the combination of text data and revision measures in explaining contemporaneous $CAR$. The DM tests confirm that report text and forecast revision contain distinct information sets that are valued by investors.
@@
The explanatory power might be derived from quantitative information in the numbers mentioned in report text. I conduct robustness checks using content with numbers removed. To address the concern, I exclude numerical data from the written reports before generating text embeddings and report model performance in Table A4. The $R^{2}_{\mathrm{OOS}}$ of the number-free text embedding input model is 10.95%, significantly higher than that of quantitative revision measures (t-statistic = 2.40). The robustness check suggests that the superior performance of text-based models is not driven by the numbers embedded in text.
@@
LLMs exhibit improved language skills as the parameter size increases. I provide a quantitative comparison of LLMs with varying numbers of parameters, including the BERT model, OpenAI text-embedding-3-small model, and LLaMA-3 model. The LLaMA models have a substantially larger number of parameters compared to the BERT and OpenAI models. Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 model. The input in Panel B is report text embeddings, comparable to column (3) in Panel A. A positive t-statistic indicates that the model outperforms the LLaMA-2 benchmark. Overall, the BERT and OpenAI models underperform the LLaMA-2 model. The LLaMA-3 model gives an $R^{2}_{\mathrm{OOS}}$ of 9.66%, which is slightly better but statistically insignificant compared to the LLaMA-2 model. The results in Table 2 suggest that the out-of-sample information content of analyst report text is fairly robust across different LLMs. Pairwise comparisons between LLMs also shed light on how the scale of the model influences performance capabilities.
To test the robustness of the information content in these contemporaneous measures, I also use $CAR$ measures in 2-day and 5-day windows, as applied in Huang et al. (2014) and Asquith et al. (2005), respectively. Additionally, I examine the abnormal stock return on the days that reports are released. In Table A5, I find a similar magnitude in $R^{2}_{\mathrm{OOS}}$ for explaining $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% for the out-of-sample period from 2015 to 2023.
@@
In Table A6, I present the information content of analyst report text across the Fama-French 12 industries. The $R^{2}_{\mathrm{OOS}}$ exceeds 10% in five industries: Shops, Other, Manufacturing, Chemicals, and Durables. Notably, the $R^{2}_{\mathrm{OOS}}$ and DM t-statistics are significantly lower in the Non-Durables, Telecommunications, Energy, and Utilities sectors. The industries with higher $R^{2}_{\mathrm{OOS}}$ typically involve complex products and services that benefit from detailed analyst insights, whereas those with lower $R^{2}_{\mathrm{OOS}}$ may operate in more stable or regulated environments where analyst reports add less incremental value. In Figure A3, I plot analyst information value by Fama-French 12 industries, revealing a pattern consistent with the information content results.
To assess the economic magnitude, I perform OLS regressions specified as follows:
@@
This model incorporates the out-of-sample estimation of $CAR_{[-1,+1]}$ using report text embeddings or forecast revisions as input. For instance, observations in 2015 are fitted using a Ridge regression model trained on samples from 2000 to 2014. The fitted value of $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension. For comparison, I estimate $\widehat{CAR}_{rev}$ using the three forecast revision measures, following the same procedure.
Table 3 summarizes the summary statistics of the sample used and the regression results. The out-of-sample period starts in 2015. I constrain the samples to include only those with valid previous and current stock recommendations, earnings forecasts, and target prices. The final sample for estimating Equation (10) contains 28,837 observations.
Column (1) of Panel B demonstrates that quantitative measures of recommendation revisions, earnings forecast revisions, and target price revisions all contribute to explaining variations in contemporaneous abnormal stock returns, corroborating the findings of Brav and Lehavy (2003) and Huang et al. (2014). Column (2) reveals that $\widehat{CAR}_{txt}$ alone yields an $R^{2}$ of 10.5%, closely aligning with the $R^{2}_{\mathrm{OOS}}$ value in Table 2. The inclusion of both analyst revision measures and $\widehat{CAR}_{txt}$ as regressors elevates $R^{2}$ from 9.1% in column (1) to 14.9% in column (3), underscoring the incremental value of text information.
The positive and significant coefficient on $\widehat{CAR}_{txt}$ indicates that the information in the report text is not subsumed by forecast revisions. Economically, a one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness increases the 3-day abnormal return by 10 basis points, 40 basis points, 90 basis points, and 120 basis points, respectively.
In columns (4) and (5) of Panel B, I report the coefficient estimates for $\widehat{CAR}_{rev}$, which combines the information from all revision metrics. The $\widehat{CAR}_{rev}$ model alone achieves an $R^{2}$ of 8.9%, while combined input improves it to 14.7%. Economically, a one standard deviation increase in $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{txt}$ corresponds to increases of 100 and 120 basis points in realized $CAR$, respectively. These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
@@
I further control for $SUE$ and $Prior\_CAR$ in columns (6)-(7) to capture recent news and earnings surprises. The coefficient estimates on $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{rev}$ remain unchanged. The stability of the coefficients suggests that market-valued analyst information is distinct from recent earnings news and is not simply a reaction to past returns.
3.2.2 Information Content around Earnings Announcements
Some studies have emphasized the relationship between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), in terms of how these two information sources reinforce each other in sequence. I examine the information content of sell-side analyst reports around earnings announcements, the most significant corporate events.
@@ 3.2.2 Information Content around Earnings Announcements
Following Chen et al. (2010), I group analyst reports into 13 weekly bins surrounding earnings announcement dates and aggregate their information content by week. Panel A in Table 4 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for each week. Consistent with Huang et al. (2018), a large portion of analyst reports are released immediately after earnings announcements. Notably, the information content is significant only in the first week, with an $R^{2}_{\mathrm{OOS}}$ of approximately 10% and a t-stat of 10.17. This performance drops sharply in the second week, turns negative in the third and fourth weeks.
@@
The distinct performance observed across weekly bins can be attributed to Ridge models effectively capturing dominant patterns in analyst reports surrounding earnings announcements. To account for potential temporal variation, I train Ridge models using separate samples of analyst reports released surrounding and beyond earnings announcements, enabling the models to extract distinct information content based on the timing of the reports. Panel B in Table 4 compares the model performance during earnings announcement periods and non-earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows. The $R^{2}_{\mathrm{OOS}}$ of models trained during earnings announcement periods is consistently higher than those trained beyond these periods. The $R^{2}_{\mathrm{OOS}}$ and t-statistics decline as the windows lengthen, declining from a peak of 11.84% to 7.29% over the course of the earnings announcement week. This trend suggests that the market values information from analysts who react swiftly to earnings announcements.
@@
One potential concern is that analysts may simply restate information from earnings calls, and that market reactions are driven by the earnings announcements themselves rather than analyst reports. If analysts merely restate this information, the measured information content of their reports might actually reflect the text from earnings conference calls. To address this issue, I use earnings conference call transcripts to differentiate the information content between the two text sources. I analyze reports released within one day after earnings announcements, when analysts are most likely to simply restate earnings information.
@@
In column (1) of Panel C, I present the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using embeddings from earnings conference call transcript embeddings as input. [^10] Given the extensive length of earnings conference call transcripts (median length of 13,029 tokens), I employed the LLaMA-3-8B model for generating text embeddings in Panel C. This model accommodates up to 8,192 tokens, compared to the 4,096 token limit of the LLaMA-2-13B model used elsewhere in our study. I truncate the text at the length limits to maintain maximum information. The $R^{2}_{\mathrm{OOS}}$ is 4.20% for the S&P 100 firms, with a DM t-stat of 5.16. Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods. The input comprised the mean of all report text embeddings on the day following the earnings announcement. The Ridge regression yielded an $R^{2}_{\mathrm{OOS}}$ of 9.72%, as shown in column (2), notably surpassing the performance of conference call transcripts alone.
@@
Incorporating both analyst report embeddings and earnings conference call embeddings in Ridge regressions resulted in a significant improvement, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% in column (3). The comparison t-statistic between the transcript-only model and the combined model was 5.12, demonstrating substantial incremental value provided by analyst reports. These findings collectively refute the idea that analysts merely restate information from earnings announcements. Instead, they emphasize the unique and timely value analysts bring to the market.
3.2.3 What Content in Analyst Reports is Valued by Market Investors?
After showing that information in report text is significant, a key question arises: what specific content in analyst reports is the most valuable to the market? To systematically evaluate topic importance, I employ the Shapley value approach to attribute the text-based $R^{2}_{\mathrm{OOS}}$ to the 17 topics described in Section 2.1.1. I use sentence-segmented embeddings as report representation. The sentence-segmented approach mitigates the concern of contextual information from one topic influencing the embeddings of tokens in another. The $R^{2}_{\mathrm{OOS}}$ of text input decreases by 2.25% when I alternate from full-context embeddings to sentence-segmented embeddings. This reduction roughly quantifies the magnitude of contextual information content across sentences.
Intuitively, Shapley values are derived by calculating the reduction in $R^{2}_{\mathrm{OOS}}$ when excluding a specific topic’s embedding across all possible topic combinations. These reductions are then averaged over the possible different permutations into a single importance measure. By design, the sum of Shapley values across all 17 topics precisely equals the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings, ensuring an efficient attribution of explanatory power.

### v3

@@ 3.2 Information Content of Analyst Reports
This section investigates market reactions to analysts’ quantitative and qualitative outputs. I employ the Shapley value decomposition methodology introduced in Section 2.4 to attribute the information content across distinct topics.
3.2.1 Quantitative and Qualitative Information
@@
Table 3 compares the information content embedded in numerical and textual measures. The analysis employs Ridge regression as specified in Equations 1 and 2, with ${CAR}_{[-1,+1]}$ centered at the report release date serving as the dependent variable. For reports released on non-trading days, I adjust t=0 to the next trading day recorded in the CRSP database. The estimation utilizes distinct input configurations with an expanding window approach, where training sets span from 2000 to the year preceding each target year.
@@
Panel A presents the performance of four input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
@@
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for forecast revision input, consistent with existing literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst revisions. The incorporation of additional 14 numerical features in column (2) yields a modest improvement to 9.08% $R^{2}_{\mathrm{OOS}}$. Both revision and numerical input models significantly outperform the zero benchmark, as evidenced by the Diebold-Mariano test t-statistics reported in columns (2) and (4).
@@
Columns (5) and (6) reveal that Ridge models with text embeddings achieve an $R^{2}_{\mathrm{OOS}}$ of 10.19% with a t-statistic of 8.20, highlighting the substantial information content of report text. The $R^{2}_{\mathrm{OOS}}$ remains stable from 2015 to 2023, except during the 2020 pandemic recession. A DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, indicating substantial information content of report text relative to forecast revisions.
@@
Addressing the look-ahead bias concern raised by Sarkar and Vafa (2024) regarding pre-trained LLMs, I examine the $R^{2}_{\mathrm{OOS}}$ in 2023. Since the employed LLMs were trained on pre-2023 data, this period represents true out-of-sample performance, validating the robustness of the main findings.
@@
Columns (7) and (8) examine the model performance when integrating both forecast revisions and text embeddings. This combined approach achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, representing improvements of 3.27% and 2.09% over revision-only and text-only models, respectively. The DM t-statistics of 3.95 and 3.77 for comparisons between combined input versus revision input and text input demonstrate that report text and forecast revisions contain distinct, complementary information sets valued by investors.
@@
The explanatory power might stem from quantitative information embedded in the numbers mentioned within the text. To address this concern, I conduct robustness checks using content with numbers removed. Table A4 presents the performance of models using number-free text embeddings. The resulting $R^{2}_{\mathrm{OOS}}$ of 10.95% significantly exceeds that of quantitative revision measures (t-statistic = 2.40), suggesting that the superior performance of text-based models is not driven by embedded numerical content.
@@
I examine the scaling law relationship by comparing LLMs with varying parameters, including BERT, OpenAI text-embedding-3-small, and LLaMA-3 models. The LLaMA models feature significantly more parameters than the BERT and OpenAI models. Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 benchmark. The input in Panel B is report text embeddings, comparable to column (3) in Panel A. A positive t-statistic indicates superior performance relative to the LLaMA-2 model. While smaller models like BERT and OpenAI underperform LLaMA-2, the larger LLaMA-3 model shows only marginally better performance with an $R^{2}_{\mathrm{OOS}}$ of 9.66%, though the improvement is statistically insignificant. These results demonstrate the robustness of analyst report text information content across different LLMs and illuminate the relationship between model scale and performance capabilities.
To validate the robustness of contemporaneous measures, I examine alternative $CAR$ windows, including 2-day and 5-day intervals following Huang et al. (2014) and Asquith et al. (2005), respectively. Table A5 shows consistent $R^{2}_{\mathrm{OOS}}$ magnitudes for $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% during the 2015-2023 out-of-sample period.
@@
Table A6 presents the cross-industry analysis of analyst report text information content across Fama-French 12 classifications. Five industries exhibit $R^{2}_{\mathrm{OOS}}$ exceeding 10%: Shops, Other, Manufacturing, Chemicals, and Durables. Conversely, Non-Durables, Telecommunications, Energy, and Utilities sectors show significantly lower $R^{2}_{\mathrm{OOS}}$ and DM t-statistics. This pattern, visualized in Figure A3, suggests that industries with complex products and services derive greater value from detailed analyst insights, while those operating in stable or regulated environments benefit less from incremental analyst information.
To evaluate the economic significance, I estimate OLS regressions specified as follows:
@@
This model incorporates out-of-sample $CAR_{[-1,+1]}$ estimations using either report text embeddings or forecast revisions as input. The estimates for 2015, for example, are derived from Ridge regression models trained on 2000-2014 data. The $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension, while $\widehat{CAR}_{rev}$ is estimated using three forecast revision measures under identical procedures.
Table 4 presents sample summary statistics and regression results. The analysis sample is restricted to observations with valid current and previous stock recommendations, earnings forecasts, and target prices. Panel B’s Column (1) confirms that recommendation revisions, earnings forecast revisions, and target price revisions all contribute significantly to explaining contemporaneous abnormal stock return variations, consistent with Brav and Lehavy (2003) and Huang et al. (2014). Column (2) shows that $\widehat{CAR}{txt}$ alone achieves an $R^{2}$ of 10.5%. When both analyst revision measures and $\widehat{CAR}{txt}$ are included as regressors in column (3), the $R^{2}$ increases to 14.9%. In economic terms, one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness corresponds to 3-day abnormal return increases of 10, 40, 90, and 120 basis points, respectively. Columns (4) and (5) examine $\widehat{CAR}{rev}$, which aggregates all revision metrics. The $\widehat{CAR}{rev}$-only model achieves an $R^{2}$ of 8.9%, while combined inputs reach 14.7%. These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
@@
Columns (6)-(7) introduce $SUE$ and $Prior_{C}AR$ controls to capture earnings surprises and recent news. The stability of $\widehat{CAR}{rev}$ and $\widehat{CAR}{txt}$ coefficients suggests that market-valued analyst information is distinct from recent earnings news and does not merely reflect reactions to past returns.
3.2.2 Earnings Announcements
Building on research examining the interplay between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), I analyze the information content of sell-side analyst reports surrounding earnings announcements.
@@ 3.2.2 Earnings Announcements
Following Chen et al. (2010), I group analyst reports into 13 weekly bins around earnings announcement dates. Panel A in Table 5 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for weekly report samples. Consistent with Huang et al. (2018), analyst reports cluster predominantly in the immediate post-earnings announcement period. The information content only exhibits strong significance only in the first week, achieving an $R^{2}_{\mathrm{OOS}}$ of approximately 10% with a t-statistic of 10.17.
@@
The variation in performance across weekly bins may reflect Ridge models capturing dominant patterns in analyst reports issued around earnings announcements. To address potential temporal variation, I train separate Ridge models on analyst reports released during earnings announcement periods and those issued beyond these periods, allowing the models to extract distinct information content based on report timing. Panel B in Table 5 compares the model performance within and beyond earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows. Models trained on earnings announcement periods consistently outperform those trained on non-announcement periods. The $R^{2}_{\mathrm{OOS}}$ and t-statistics exhibit a declining pattern as windows lengthen, decreasing from 11.84% to 7.29% over the earnings announcement week, suggesting markets particularly value swift analyst responses to earnings news.
@@
To address concerns that analysts merely restate earnings call information and that market reactions primarily reflect earnings announcements rather than analyst insights, I conduct a comparative analysis of information content between these two sources. I focus on reports released one day after earnings announcements, when analysts would be most likely to simply repackage earnings information.
@@
Column (1) of Panel C presents the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using earnings conference call transcript embeddings as input. [^10] Due to the length of earnings conference call transcripts (median 13,029 tokens), I employ the LLaMA-3-8B model for text embeddings in Panel C, which accommodates 8,192 tokens versus the 4,096-token limit of the LLaMA-2-13B model used elsewhere. Texts are truncated at these limits to maximize information retention. The $R^{2}_{\mathrm{OOS}}$ reaches 4.20% with a DM t-statistic of 5.16. Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods. Using report text embeddings from the day following earnings announcements, this analysis achieves an $R^{2}_{\mathrm{OOS}}$ of 9.72% (column 2), substantially outperforming conference call transcripts alone.
@@
The combination of analyst report and earnings conference call embeddings yields further improvements, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% (column 3). A comparison t-statistic of 5.12 between transcript-only and combined models demonstrates analysts’ substantial incremental contribution. These findings collectively refute the notion that analysts merely restate earnings information, instead highlighting their distinct and timely value to the market.
3.2.3 What Content in Analyst Reports is Valued?
Having established the significance of report text information, I examine which specific content components most influence market response. I employ the Shapley value methodology to decompose the text-based $R^{2}_{\mathrm{OOS}}$ across the 17 topics detailed in Table LABEL:tab:_topic. The analysis uses sentence-segmented embeddings as report representations to mitigate cross-topic contextual effects. When switching from full-context to sentence-segmented embeddings, the $R^{2}_{\mathrm{OOS}}$ decreases by 2.25%, quantifying the magnitude of inter-sentence contextual information.